OpenStudy (anonymous):

Calculate derivative of \[f(x)=\log(2+sinx)\]

geerky42 (geerky42):

Perhaps it's best if you explain how you got that answer?

OpenStudy (anonymous):

yes that would be nice

geerky42 (geerky42):

@mr.singh

OpenStudy (anonymous):

@geerky can you explain it pls ?

geerky42 (geerky42):

You can "see" it as composition of functions: for example: Let \(g(x)\) be \(\log(x)\) and \(h(x)\) be \(2+\sin(x)\) So \(f(x)\) would be \(g(h(x))\) So you can apply chain rule, where \(f'(x) = g'(h(x))\cdot h'(x)\) You agree that \(g'(x) = \dfrac{1}{x}\) and \(h'(x)=\cos x\), right? So \(f'(x) = g'(h(x))\cdot h'(x) = \left[\dfrac{1}{2+sinx}\right]\cdot\cos x\\~~~~~~~~~=\boxed{\dfrac{\cos x}{2+\sin x}}\)

geerky42 (geerky42):

Is that clear?

OpenStudy (anonymous):

hmm i need a little bit time to understand i think

geerky42 (geerky42):

You know chain rule, right? You basically derivative "outer," leave "inner" untouched, then you multiply it by derivative of "inner"

OpenStudy (anonymous):

yes i know it, ok now its better thank you very much geerky

geerky42 (geerky42):

Ok welcome.